# Text-to-Vision Generation with Scene Graph Programming

### v2

@@ D.3 Case study: Pairwise fine-grained model comparison
Figure 14: Stable Diffusion v2-1 vs. Stable Diffusion 3 Medium on average VQA Score in fine-grained categories.
Figure 15: PixArt-$\Sigma$ vs. Stable Diffusion 3 Medium on average VQA Score in fine-grained categories.
Figure 16: FLUX.1-schnell vs. Stable Diffusion 3 Medium on average VQA Score in fine-grained categories.
Figure 17: PixArt-$\Sigma$ vs. FLUX.1-schnell on average VQA Score in fine-grained categories.
Appendix E Details of Application 1: Self-Improving Models (Section 4)
E.1 Experiment details
E.1.1 Captions Preparation
@@
For comparative analysis, we leveraged the Conceptual Captions (CC3M) [9] dataset, a large-scale benchmark containing approximately 3.3 million image-caption pairs sourced from web alt-text descriptions. CC3M is renowned for its diverse visual content and natural language expressions, encompassing a wide range of styles, contexts, and semantic nuances.
@@
E.1.2 Dataset Construction and Selection Strategies
For the captions generated by Generate Any Scene, we employed a top-scoring selection strategy to construct the fine-tuning training dataset, using a random selection strategy as a baseline for comparison. Specifically, for each prompt, the model generated eight images. Under the top-scoring strategy, we evaluated the generated images using the VQA score and selected the highest-scoring image as the best representation of the prompt. This process yielded 10k top-ranked images per iteration, from which the top 25% (approximately 2.5k images) with the highest VQA scores were selected to form the fine-tuning dataset.
In the random selection strategy, one image was randomly chosen from the eight generated per prompt, and 25% of these 10k randomly selected images were sampled to create the fine-tuning dataset, maintaining parity in data size.
For the CC3M dataset, each prompt was uniquely paired with a real image. From the 10k real image-caption pairs sampled from CC3M, the top 25% with the highest VQA scores were selected as the fine-tuning training dataset. This ensured consistency in data size and selection criteria across all methods, facilitating a rigorous and equitable comparison of fine-tuning strategies.
E.1.3 Fine-tuning details
We fine-tuned the Stable Diffusion v1-5 using the LoRA technique. The training was conducted with a resolution of 512 $\times$ 512 for input images and a batch size of 8. Gradients were accumulated over two steps. The optimization process utilized the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, an $\epsilon$ value of $1\times 10^{-8}$, and a weight decay of $10^{-2}$. The learning rate was set to $1\times 10^{-4}$ and followed a cosine scheduler for smooth decay during training. To ensure stability, a gradient clipping threshold of 1.0 was applied. The fine-tuning process was executed for one epoch, with a maximum of 2500 training steps. For the LoRA-specific configurations, we set the rank of the low-rank adaptation layers and the scaling factor $\alpha$ to be 128.
After completing fine-tuning for each epoch, we set the LoRA weight to 0.75 and integrate it into Stable Diffusion v1-5 to guide image generation and selection for the next subset. For the CC3M dataset, images from the subsequent subset are directly selected.
@@ E.1.3 Fine-tuning details
E.2 More results of fine-tuning models
Aside from our own test set and GenAI benchmark, we also evaluated our fine-tuned Text-to-Image generation models on the Tifa Bench (Figure 18), where we observed the same trend: models fine-tuned with our prompts consistently outperformed the original Stable Diffusion v1-5 and CC3M fine-tuned models.
Figure 18: Results for Application 1: Self-Improving Models. Average TIFA score of Stable Diffusion v1-5 fine-tuned with different data over TIFA Bench.
Appendix F Details of Application 2: Distilling limitations (Section 5)
F.1 Collecting hard concepts
We selected 81 challenging object concepts where Stable Diffusion v1-5 and DaLL-E 3 exhibit the largest gap in VQA Score. To determine the score for each concept, we calculated the average VQA score of the captions containing that specific concept. The full list of hard concepts is shown below:
@@ F.1 Collecting hard concepts
F.2 Experiment details
We conducted targeted fine-tuning experiments on Stable Diffusion v1-5 to evaluate Generate Any Scene’s effectiveness in distilling model compositionality and learning hard concepts. For each task, we selected a dataset of 778 Generate Any Scene captions paired with images generated by DaLL-E 3. For compositionality, we selected multi-object captions from the existing dataset of 10k Generate Any Scene captions and paired them with the corresponding images generated by DaLL-E 3. To address hard concept learning, we first used Stable Diffusion v1-5 to generate images based on the 10k Generate Any Scene captions and identified the hard concepts with the lowest VQA scores. These concepts were then used to create a subset of objects, which we recombined into our scene-graph based captions with complexity levels ranging from 3 to 9. Finally, we used DaLL-E 3 to generate corresponding images for these newly composed captions.
The fine-tuning configurations were consistent with those used in the self-improving setup (Appendix E.1.3). To accommodate the reduced dataset size, the maximum training steps were set to 1000.
@@ F.2 Experiment details
Appendix G Details of Application 3: Generated content detector (Section 3)
G.1 Experiment details
@@
We employed ViT-T [73] and ResNet-18 [20] as backbones for the detection models. Their pretrained parameters on ImageNet-21k were frozen, and the final classification head was replaced with a linear layer using a sigmoid activation function to predict the probability of an image being AI-generated. During training, We used Binary Cross-Entropy (BCE) as the loss function, and the AdamW optimizer was applied with a learning rate of $2e^{-3}$. Training was conducted with a batch size of 256 for up to 50 epochs, with early stopping triggered after six epochs of no improvement in validation performance.
@@
For GenImage, we used validation data from four models: SD v1.4, SD v1.5, MidJourney, and VQDM. Each validation set contained 8k real images and 8k generated images. For Generate Any Scene, we sampled 10k real images from CC3M and paired them with 10k generated images from each of the following models: Stable Diffusion v2-1, PixArt-$\alpha$, Stable Diffusion 3 Medium, and Playground v2.5. This created distinct test sets for evaluating model performance across different synthetic data sources.
Detector Data Scale (2N)
@@ G.2 Results
Detector Data Scale (2N) SDv1.5 VQDM Midjourney

### v3

@@ A.4.5 Case study: Pairwise fine-grained model comparison
Appendix B Details of Taxonomy of Visual Concepts
Our metadata is further organized using a well-defined taxonomy, enhancing the ability to generate controllable captions. This hierarchical taxonomy not only facilitates the creation of diverse scene graphs, but also enables fine-grained and systematic model evaluation.
Objects.
To enhance the comprehensiveness and taxonomy of object data, we leverage noun synsets and the structure of WordNet (Miller, 1995). In WordNet, a physical object is defined as "a tangible and visible entity; an entity that can cast a shadow." Following this definition, we designate the physical object as the root node, constructing a hierarchical tree with all 28,787 hyponyms under this category as the set of objects in our model.
Attributes.
The attributes of a scene graph represent properties or characteristics associated with each object. We classify these attributes into nine primary categories. For color, we aggregate 677 unique entries sourced from Wikipedia (Wikipedia Contributors, 2024). The material category comprises 76 types, referenced from several public datasets (Vecchio & Deschaintre, 2024; Bell et al., 2015; Xue et al., 2017). The texture category includes 42 kinds from the Describable Textures Dataset (Cimpoi et al., 2014), while the architectural style encompasses 25 distinct styles (Xu et al., 2014). Additionally, we collect 85 states, 41 shapes, and 24 sizes. For human descriptors, we compile 59 terms across subcategories, including body type and height. Finally, we collect 465 common adjectives covering general characteristics of objects to enhance the descriptive richness of our scene graphs.
Relationships.
We leverage the Robin dataset (Park et al., 2025) as the foundation for relationship metadata, encompassing six key categories: spatial, functional, interactional, social, emotional, and symbolic. With 10,492 relationships, the dataset provides a comprehensive and systematic repository that supports modeling diverse and complex object interactions. Its extensive coverage captures both tangible and abstract connections, forming a robust framework for accurate scene graph representation.
Scene Attributes.
In Text-to-Vision generation tasks, people mainly focus on creating realistic images and art from a text description (Saharia et al., 2022; Ramesh et al., 2022; Betker et al., 2023). For artistic styles, we define scene attributes using 76 renowned artists, 41 genres, and 126 painting styles from WikiArt (Saleh & Elgammal, 2015), along with 29 common painting techniques. For realistic imagery, we construct camera settings attributes across 6 categories: camera models, focal lengths, perspectives, apertures, depths of field, and shot scales. The camera models are sourced from the 1000 Cameras Dataset (Crawford, 2018), while the remaining categories are constructed based on photography knowledge and common captions in Text-to-Vision generation tasks (Brooks et al., 2024; Wang et al., 2022). To control scene settings, we categorize location, weather and lighting attributes, using 430 diverse locations from Places365 (López-Cifuentes et al., 2020), alongside 76 weathers and 57 lighting conditions. For video generation, we introduce attributes that describe dynamic elements. These include 12 types of camera rig, 30 distinct camera movements, 15 video editing styles, and 27 temporal spans. The comprehensive scene attributes that we construct allow for the detailed and programmatic Text-to-Vision generation generation.
Appendix C Details of self-improving models with synthetic captions (Section 3)
C.1 Experiment details
C.1.1 Captions Preparation
@@
For comparative analysis, we leveraged the Conceptual Captions (CC3M) (Changpinyo et al., 2021) dataset, a large-scale benchmark containing approximately 3.3 million image-caption pairs sourced from web alt-text descriptions. CC3M is renowned for its diverse visual content and natural language expressions, encompassing a wide range of styles, contexts, and semantic nuances.
@@
C.1.2 Dataset Construction and Selection Strategies
For the captions generated by Generate Any Scene, we employed a top-scoring selection strategy to construct the fine-tuning training dataset, using a random selection strategy as a baseline for comparison. Specifically, for each caption, the model generated eight images. Under the top-scoring strategy, we evaluated the generated images using the VQA score and selected the highest-scoring image as the best representation of the caption. This process yielded 10K top-ranked images per iteration, from which the top 25% (approximately 2.5k images) with the highest VQA scores were selected to form the fine-tuning dataset.
In the random selection strategy, one image was randomly chosen from the eight generated per caption, and 25% of these 10K randomly selected images were sampled to create the fine-tuning dataset, maintaining parity in data size.
For the CC3M dataset, each caption was uniquely paired with a real image. From the 10K real image-caption pairs sampled from CC3M, the top 25% with the highest VQA scores were selected as the fine-tuning training dataset. This ensured consistency in data size and selection criteria across all methods, facilitating a rigorous and equitable comparison of fine-tuning strategies.
C.1.3 Fine-tuning details
We fine-tuned the SDv1.5 using the LoRA technique. The training was conducted with a resolution of 512 $\times$ 512 for input images and a batch size of 8. Gradients were accumulated over two steps. The optimization process utilized the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, an $\epsilon$ value of $1\times 10^{-8}$, and a weight decay of $10^{-2}$. The learning rate was set to $1\times 10^{-4}$ and followed a cosine scheduler for smooth decay during training. To ensure stability, a gradient clipping threshold of 1.0 was applied. The fine-tuning process was executed for one epoch, with a maximum of 2500 training steps. For the LoRA-specific configurations, we set the rank of the low-rank adaptation layers and the scaling factor $\alpha$ to be 128.
After completing fine-tuning for each epoch, we set the LoRA weight to 0.75 and integrate it into SDv1.5 to guide image generation and selection for the next subset. For the CC3M dataset, images from the subsequent subset are directly selected.
@@ C.1.3 Fine-tuning details
In Figure 15, we present results using our captions and the CC3M captions. The model fine-tuned with captions generated by Generate Any Scene demonstrates superior performance in terms of text semantic relevance and the generation of complex compositional scenes.
Figure 15: Visualization of Different Caption Fine-Tuning.
C.2 Evaluation on TIFA bench
Aside from our own test set and GenAI benchmark, we also evaluated our fine-tuned Text-to-Image generation models on the Tifa Bench (Figure 16), where we observed the same trend: models fine-tuned with our captions consistently outperformed the original SDv1.5 and CC3M fine-tuned models.
Figure 16: Results for Application 1: Self-Improving Models. Average TIFA score of SDv1.5 fine-tuned with different data over TIFA Bench.
C.3 Additional real-data baselines
Setup.
We conduct more experiments comparing Generate Any Scene synthetic captions to other real-world caption sources. We sampled 10K captions from MS-COCO-2017 and LAION-COCO for one-epoch LoRA fine-tuning under same experimental settings. The results on Generate Any Scene test set are summarized in Table 9.
Table 9: Self-improvement on Generate Any Scene Test (VQA). One-epoch finetuning, equal budget.
Method VQA $\uparrow$
Baseline (SD-1.5) 0.508
MS-COCO-2017 0.508
LAION-COCO 0.510
CC3M 0.508
GAS (Random) 0.524
GAS (Top-Score) 0.530
Findings.
Fine‑tuning with MS‑COCO‑2017 and LAION‑COCO captions yields results similar to CC3M, with none surpassing the significant improvements achieved by our Generate Any Scene captions. We think that although MS-COCO-2017 and LAION captions are generally high-quality and well‑aligned with images, they offer limited compositional diversity. These additional results confirm that the observed gains are not specific to CC3M but generalize across other widely used real-caption datasets. This further supports our claim that the compositional diversity of Generate Any Scene synthetic captions drives the improvement.
C.4 Full fine-tuning vs. LoRA fine-tuning
Setup.
We replicate the self-improvement pipeline with full fine-tuning and compare three strategies: Generate Any Scene captions with high‑score selection, Generate Any Scene captions with random selection, and CC3M captions as the real‑data baseline. The results are shown in Tables 10 and 11.
Table 10: Results on Generate Any Scene test set under full fine-tuning. (VQA Score)
Method Iter-1 Iter-2 Iter-3
Baseline 0.508
CC3M (Full FT) 0.496 0.518 0.519
GAS (Rand, Full FT) 0.510 0.519 0.520
GAS (Top, Full FT) 0.510 0.534 0.540
Table 11: Results on GenAI-Bench under full fine-tuning. (VQA Score)
Method Iter-1 Iter-2 Iter-3
Baseline 0.617
CC3M (Full FT) 0.589 0.619 0.622
GAS (Rand, Full FT) 0.599 0.621 0.617
GAS (Top, Full FT) 0.620 0.626 0.634
Findings.
Using our Generate Any Scene captions with high score selection not only improves performance consistently across iterations but also surpasses CC3M at every stage. The full fine-tuning results confirm that our captions and strategy’s effectiveness is not dependent on the specific training approach (LoRA vs. full fine-tuning). The consistent improvement patterns across both evaluation benchmarks validate the robustness of our iterative self-improvement framework.
Appendix D Details of distilling targeted capabilities (Section 4)
D.1 Collecting hard concepts
We selected 81 challenging object concepts where SDv1.5 and DaLL-E 3 exhibit the largest gap in VQA Score. To determine the score for each concept, we calculated the average VQA score of the captions containing that specific concept. The full list of hard concepts is shown below:
@@ D.1 Collecting hard concepts
D.2 Experiment details
We conducted targeted fine-tuning experiments on SDv1.5 to evaluate Generate Any Scene’s effectiveness in distilling model compositionality and learning hard concepts. For each task, we selected a dataset of 778 Generate Any Scene captions paired with images generated by DaLL-E 3. For compositionality, we selected multi-object captions from the existing dataset of 10K Generate Any Scene captions and paired them with the corresponding images generated by DaLL-E 3. To address hard concept learning, we first used SDv1.5 to generate images based on the 10K Generate Any Scene captions and identified the hard concepts with the lowest VQA scores. These concepts were then used to create a subset of objects, which we recombined into our scene-graph based captions with complexity levels ranging from 3 to 9. Finally, we used DaLL-E 3 to generate corresponding images for these newly composed captions.
The fine-tuning configurations were consistent with those used in the self-improving setup (Appendix C.1.3). To accommodate the reduced dataset size, the maximum training steps were set to 1000.
@@ D.2 Experiment details
D.3 Benchmark against web-crawled caption–image pairs
Setup.
We conduct additional experiments to benchmark against alternative data sourcing strategies, specifically comparing our DaLL-E 3 distillation approach with web-scraped real images. Using the Bing Image Search API, we retrieve images matching our multi‑object and hard‑concept captions and constructed two datasets of equivalent scale for comparison. We then apply the same fine‑tuning setup described in Application 2. The results are shown in Table 12:
Table 12: Comparison of VQA scores from targeted fine-tuning on different data sources. (SDv1.5)
Test Set Original DaLL-E 3 Distill Web-crawled
Hard Concept 0.303 0.361 0.258
Multi-object 0.271 0.325 0.264
Findings.
The results show that web-scraped images not only failed to improve performance but actually degraded model capabilities.
Upon examination of the retrieved images, we identify several critical issues. The web-crawled images contain significant noise, including watermarks, overlaid text, and irrelevant visual element. Our hard concept and multi-object captions feature high compositional complexity and novel object combinations that rarely exist in real-world photographs. The retrieved images show poor relevance to our systematically designed compositional scenarios, as real-world images cannot adequately represent the diverse and controlled compositional variations we programmatically generate. Thus, training on such misaligned data appears to introduce incorrect visual-textual associations, leading to performance degradation rather than improvement.
Table 13: VQA Score of targeted distillation on FLUX.1-dev.
Test Set Original Fine-tuned
Hard Concept 0.303 0.361
Multi-object 0.271 0.325
D.4 Distillation on Flux.1-dev
Setup.
We further apply our distillation framework to FLUX.1-dev, a current SOTA open-source model, using DaLL-E 3 ‑generated images of hard concepts and multi‑object captions to distill these capabilities into FLUX.1-dev. The results are shown in the Table 13:
Findings.
The results demonstrate that our approach’s effectiveness extends to state-of-the-art models (FLUX.1-dev). The distillation approach yields substantial improvements on challenging compositional tasks.
Appendix E Details of reinforcement learning with a synthetic reward function (Section 5)
E.1 Training data preparation
We adopt SimpleAR-0.5B-SFT (Wang et al., 2025) as our base model. Given that SImpleAR-0.5B-SFT is pretrained on high-quality real image datasets such as LAION (Schuhmann et al., 2022) and CC3M (Sharma et al., 2018), we aim to mitigate potential distributional shift between the original training data and the reinforcement learning phase. To this end, we perform metadata pre-selection for Generate Any Scene by analyzing the frequency of each object category appearing in the LAION dataset. Leveraging the controllable compositional capabilities of Generate Any Scene, we filter object categories by selecting the top 10% most frequent entries and constrain scene complexity to 3–6 objects per scene. Based on these conditions, we synthesize a set of 10K captions, ensuring semantic alignment with the base model’s pretraining distribution while maintaining structural and content diversity.
E.2 Experiment details
The detailed training configuration is provided in Table 14. We utilize 8 $\times$ NVIDIA H100 GPUs (80GB HBM3), with one GPU allocated for online generation using vLLM. The total training time is approximately 14 hours.
Table 14: Scene-graph based GRPO Fine-tuning Configuration for SimpleAR
Component Details
Model Name SimpleAR-0.5B-SFT
Model Size $\sim$0.5B parameters
Training Policy GRPO
Inference Engine vLLM (GPU utilization = 0.7)
Completion Length 4096 tokens
Training Epochs 1
Batch Size per Device 4
Learning Rate $1\times 10^{-5}$
Scheduler Cosine Annealing (min lr rate = 0.1)
Warm-up Ratio 0.1
Gradient Accumulation 1
Figure 17: Reward progression during scene-graph based GRPO training.
Figure 17 illustrates the reward progression during training. A noticeable improvement in reward is observed following the application of a learning rate of 1e-5 combined with a warm-up strategy. Overall, the reward increases by approximately 0.2, indicating effective learning under the adjusted training configuration.
In Table 2, we observe that the reproduced results of baseline models on DPG-Bench and GenEval Bench are slightly lower than those reported in the original paper. Considering the inherent stochasticity in generative model outputs, we cite the original results for comparison. For GenAI-Bench, all reported results are based on our own experimental evaluations.
E.3 Reward variants and ablations
Setup.
To verify the observed gains arise specifically from the scene-graph–generated QA reward, rather than simply from using any QA-based reward, we conduct experiments incorporating manually annotated QA datasets, VQAv2, as additional reward signals under the same RLHF framework. We sample 10K images from VQAv2, with corresponding QA pairs, matched them to COCO2017 captions, and apply same training frameworks to SimpleAR-0.5B-SFT with RL training. The results on GenAI Bench are shown in the table:
Table 15: GenAI Bench performance (VQA) under RLHF with different reward sources. All models start from SimpleAR-0.5B-SFT.
Method Basic $\uparrow$ Advanced $\uparrow$ All $\uparrow$
SimpleAR-0.5B-SFT 0.74 0.60 0.66
SimpleAR-0.5B-RL (CLIP) 0.75 0.60 0.67
SimpleAR-0.5B-RL (VQAv2) 0.73 0.59 0.66
SimpleAR-0.5B-RL (Ours) 0.75 0.61 0.68
Findings.
The results show that using VQAv2 captions and QA pairs as rewards yields even lower performance than CLIP‑based RL training. Furthermore, we observe minimal reward improvement from VQA signals throughout training. We attribute this to the fact that, although VQAv2 QA pairs are rich, the underlying image captions fail to cover enough visual elements, leading to a mismatch between QA pairs and captions that undermines RLHF reward alignment.
This highlights the inherent difficulty and cost of constructing high-quality image-caption and QA annotations, whereas our method leverages scene-graph structures to systematically generate synthetic caption-QA pairs at minimal cost with unique advantages.
Appendix F Details of improving generated-content detection (Section 6)
F.1 Experiment details
@@
We employed ViT-T (Wu et al., 2022) and ResNet-18 (He et al., 2016) as backbones for the detection models. Their pretrained parameters on ImageNet-21K were frozen, and the final classification head was replaced with a linear layer using a sigmoid activation function to predict the probability of an image being AI-generated. During training, We used Binary Cross-Entropy (BCE) as the loss function, and the AdamW optimizer was applied with a learning rate of $2e^{-3}$. Training was conducted with a batch size of 256 for up to 50 epochs, with early stopping triggered after six epochs of no improvement in validation performance.
@@
For GenImage, we used validation data from four models: SD v1.4, SD v1.5, MidJourney, and VQDM. Each validation set contained 8K real images and 8k generated images. For Generate Any Scene, we sampled 10K real images from CC3M and paired them with 10K generated images from each of the following models: SDv2.1, PixArt-$\alpha$, SD3 Medium, and Playground v2.5. This created distinct test sets for evaluating model performance across different synthetic data sources.
Detector
Data Scale
@@ F.2 Results
Detector
Data Scale
SDv1.5 VQDM Midjourney
@@
Appendix G Limitation
Programmatically generated prompts can be unrealistic and biased.
Linguistic diversity of programmatic prompts is limited.
Toward curriculum-aware GRPO training.
Our proposed Generate Any Scene framework plays a central role in GRPO training by providing structured scene graphs that serve as the foundation for a semantically grounded and controllable reward function. This design enables effective optimization by aligning generation objectives with fine-grained visual semantics. Beyond this, we also observe that Generate Any Scene also offers broader potential: the scene graphs it produces vary in complexity, such as in the number of objects, attributes, relationships and graph degree. These variations naturally correspond to different levels of generation difficulty and reward variance. This property suggests an opportunity for curriculum-based training, where the model could be progressively exposed to increasingly complex scene graphs. Such a strategy may improve training stability and efficiency, especially in the early stages of learning. We identify this as a promising direction for future work, further leveraging the controllability of Generate Any Scene to guide structured policy learning.